OpenStudy (anonymous):

solve each equation system by the addition method 2x+3y=23 3x+y=17

OpenStudy (anonymous):

2x+3y=23 (3x+y=17) times 3 9x + 3y = 51

OpenStudy (anonymous):

add 9x + 3y = 51 2x + 3y = 23

OpenStudy (anonymous):

i.e 9x - 2x = 7x 3y - 3y = 0 51 - 23 = 28

OpenStudy (anonymous):

therefore 7x = 28

OpenStudy (anonymous):

x = 28/7

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

9x + 3y = 51 put x = 4 9(4) + 3y = 51

OpenStudy (anonymous):

36 + 3y = 51

OpenStudy (anonymous):

3y = 51 - 36

OpenStudy (anonymous):

3y = 15

OpenStudy (anonymous):

y = 15/3

OpenStudy (anonymous):

y = 5

OpenStudy (anonymous):

Order pair is (4,5)